OpenStudy (anonymous):

Please help and i'll medal =) If 3xm+2ym-2yn-3xn=0 and "m" does not equal "n," then what is the value of y in terms of x?

OpenStudy (anonymous):

\[3xm+2ym-2yn-3xn=0\] and your job it so solve for \(y\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok leave everything with a y on the left, everything else on the right \[2ym-2yn=3xn-3xm\]

OpenStudy (anonymous):

factor out the y on the left \[y(2m-2n)=3xn-3xm\]

OpenStudy (anonymous):

ohhhh okay

OpenStudy (anonymous):

then divide to finish with \[y=\frac{3xn-3xm}{2m-2n}\]

OpenStudy (anonymous):

what about those 2 x's

OpenStudy (anonymous):

and m cannot equal n

OpenStudy (anonymous):

i hope i got the m and n straight

OpenStudy (anonymous):

the m and the n has to disappear

OpenStudy (anonymous):

the reason they wrote that m cannot equal n is because of your answer if m was equal to n the denominator would be zero and you cannot divide by zero they just wrote it to make sure the answer makes sense

OpenStudy (anonymous):

oh ic

OpenStudy (anonymous):

ok we can make them go away i think, hold on

OpenStudy (anonymous):

\[y=\frac{3xn-3xm}{2m-2n}\] factor again as \[y=\frac{3x(n-m)}{3(m-n)}\]

OpenStudy (anonymous):

typo there should be a 2 in the denominator

OpenStudy (anonymous):

bottom is 2 but yea so do they cancel out that way?

OpenStudy (anonymous):

\[y=\frac{3x(n-m)}{2(m-n)}\]

OpenStudy (anonymous):

shouldnt this be a negative

OpenStudy (anonymous):

they do cancel, but they don't leave a 1 they cancel to get -1 because \[\frac{n-m}{m-n}=-1\]

OpenStudy (anonymous):

final answer if i did not mess up with the m and n is \[y=-\frac{3}{2}x\]

OpenStudy (anonymous):

hmmm kinda confusing

OpenStudy (anonymous):

yeah i agree

OpenStudy (anonymous):

tryna figure out why the m and n's went to the left side

OpenStudy (anonymous):

oh because you wanted so solve for \(y\) so all the terms with y have to be on one side of the equal sign, everything else on the other

OpenStudy (anonymous):

hmmmm okay i got it im gonna have to go over this a couple more times but thanks satellite

OpenStudy (anonymous):

yw we can do it again slowly if you like, starting with \[3xm+2ym-2yn-3xn=0\]

OpenStudy (anonymous):

yea thatd be great

OpenStudy (anonymous):

ok it is clear you want an answer that looks like \(y=something\) right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[3xm+2ym-2yn-3xn=0\] the middle two terms have a y in them , the first and last terms do not

OpenStudy (anonymous):

yea so faactor it out right

OpenStudy (anonymous):

we leave the middle two terms on the left, and subtract \(3xm\) from both sides, also add \(3xn\) to both sides

OpenStudy (anonymous):

that is how i got to the line \[2ym-2yn=3xn-3xm\] ok to here?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now factor out the y on the left, so you can see what do divide both sides by \[y(2m-2n)=3xn-3xm\]

OpenStudy (anonymous):

now divide both sides by \(2m-2n\) and get \[y=\frac{3xn-3xm}{2m-2n}\]

OpenStudy (anonymous):

hey can i try something else

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what about m(3x+2y)-n(2y+3x=0

OpenStudy (anonymous):

but i dont understand why it becomes just (3x+2y) (m-n)=0 where does that other (3x+2y) go?

OpenStudy (anonymous):

that is right, but i am not sure how it is going to get you to \(y=expression\)

OpenStudy (anonymous):

you really need to have all the y on one side, everything else on the other

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

this is really confusing i am screwed if i get this on the GRE

OpenStudy (anonymous):

once we get to \[y=\frac{3xn-3xm}{2m-2n}\] we have achieved that goal now it is a matter of factoring and cancelling on the right

OpenStudy (anonymous):

factor a \(3x\) from the numerator and a \(2\) from the denominator, get \[y=\frac{3x(n-m)}{2(m-n)}\]

OpenStudy (anonymous):

ok im with you

OpenStudy (anonymous):

then cancel, get \[y=-\frac{3x}{2}\]

OpenStudy (anonymous):

the gimmick here is to recognize that \[\frac{n-m}{m-n}=-1\]

OpenStudy (anonymous):

so you treat the (n-m) as a variable and you divide from each other and cancel out right?

OpenStudy (anonymous):

yeah, it is always the case that \[\frac{a-b}{b-a}=-1\] since the denominator is the negative of the numerator and vice versa

OpenStudy (anonymous):

ohhh..... is that rule?

OpenStudy (anonymous):

it is just a fact

OpenStudy (anonymous):

i did not know this....

OpenStudy (anonymous):

\[\frac{5-2}{2-5}=\frac{3}{-3}=-1\] for example you can try it with your own numbers, see that it works

OpenStudy (anonymous):

hmmmmm okay i think i got it now...

OpenStudy (anonymous):

ohhhhhh.... right omfg that blew my mind

OpenStudy (anonymous):

thank so much satellite

OpenStudy (anonymous):

yw, good luck with the gre

OpenStudy (anonymous):

thanks!